# VerIde ECG Biometrics: Verification and Identification

### v1

@@ Abstract
This work studies electrocardiogram (ECG) biometrics at large scale, evaluating how strongly an ECG can be linked to an individual and, consequently, how its anonymization may be compromised.
We show that identity information is already present in tabular representations (fiducial features): even a simple MLP-based embedding network yields non-trivial performance, indicating that anonymization based solely on releasing features does not guarantee privacy.
We then adopt embedding-based deep learning models (ArcFace), first on features and then on ECG waveforms, showing a performance jump when moving from tabular inputs to waveforms, and a further gain with larger training sets and consistent normalization across train/val/test.
@@
Overall, the results show that ECG carries a measurable individual signature: re-identification is already possible with tabular features and is further amplified by embedding-based models, making privacy implications and realistic operational protocols essential to consider.
@@ 1 Introduction
The literature highlights two recurring critical issues that limit comparability across studies and the operational interpretation of results: (i) the need for clear and comparable experimental protocols (dataset choice, split criteria, session handling, acquisition conditions); and (ii) the adoption of metrics consistent with operational biometrics rather than purely “laboratory” settings, capable of describing system behavior at realistic security thresholds [22, 9, 11]. On this basis, this study analyzes ECG biometrics at large scale, using a test set of about 20,000 identities and reporting metrics and protocols that make performance interpretable in real-world applications.
In recent years, the adoption of deep learning models and embedding representations has accelerated ECG biometric systems, but heterogeneous experimental settings (dataset size, preprocessing, multi-session, and metric choice) still make direct comparison difficult. In this context, the contribution of this work is primarily methodological and experimental: rather than proposing a new loss or architecture, we aim to build and evaluate a reproducible, scalable, and operationally oriented pipeline.

### v2

# ASUGI - Azienda Sanitaria Universitaria Giuliano-Isontina

@@ Abstract
This work studies electrocardiogram (ECG) biometrics at large scale, directly addressing a critical gap in the literature: the scarcity of large-scale evaluations with operational metrics and protocols that enable meaningful standardization and comparison across studies.
We show that identity information is already present in tabular representations (fiducial features): even a simple MLP-based embedding network yields non-trivial performance, establishing a strong baseline before waveform modeling.
We then adopt embedding-based deep learning models (ArcFace), first on features and then on ECG waveforms, showing a clear performance jump when moving from tabular inputs to waveforms, and a further gain with larger training sets and consistent normalization across train/val/test.
@@
Overall, the results show that ECG carries a measurable individual signature and that large-scale testing is essential to obtain realistic, comparable metrics. The study provides an operationally grounded benchmark that helps standardize evaluation across protocols.
@@ 1 Introduction
The literature highlights two recurring critical issues that limit comparability across studies and the operational interpretation of results: (i) the need for clear and comparable experimental protocols (dataset choice, split criteria, session handling, acquisition conditions); and (ii) the adoption of metrics consistent with operational biometrics rather than purely “laboratory” settings, capable of describing system behavior at realistic security thresholds [22, 9, 11]. A third, practical gap is the scarcity of large-scale evaluations, which prevents meaningful standardization of metrics and protocols. On this basis, this study analyzes ECG biometrics at large scale, using a test set of about 20,000 identities and reporting metrics and protocols that make performance interpretable in real-world applications.
In recent years, the adoption of deep learning models and embedding representations has accelerated ECG biometric systems, but heterogeneous experimental settings (dataset size, preprocessing, multi-session, and metric choice) still make direct comparison difficult. In this context, the contribution of this work is primarily methodological and experimental: rather than proposing a new loss or architecture, we provide large-scale evidence with operational metrics to promote standardization and to build and evaluate a reproducible, scalable, and deployment-oriented pipeline.